OpenStudy (anonymous):

Write an equation in Slope-intercept form for the line that is perpendicular to the line y = 2x - 5 that passes through the given point (4,-2). Show your work.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

So slope intercept form is when you give the slope and y-intercept of the line. For this question it would be easier to start with point-slope form and then change it into slope-intercept. First find the slope.

OpenStudy (anonymous):

The slope of the original line is 2. The slope of a line that is perpendicular to another line is always the negative reciprocal of the slope of the first line. In this case it is -1/2

OpenStudy (anonymous):

Now you can just plug in information. Point slope form is (y-y1)=m(x-x1) where (x1,y1) is a point on the line and m is the slope of the line. In this question, m= -1/2 and (x1,y1) = (4,-2). So the equation is (y+2)=(-1/2)(x-4)

OpenStudy (anonymous):

Expanding this, you get y= (-1/2x)

OpenStudy (anonymous):

you think you could help me on another question?